OpenStudy (anonymous):

If 34.2 grams of lithium react with excess water, how many liters of hydrogen gas can be produced at 299 Kelvin and 1.21 atmospheres? Show all of the work used to solve this problem. 2 Li (s) + 2 H2O (l) yields 2 LiOH (aq) + H2 (g)

OpenStudy (anonymous):

mol Li = 34.2g / 6.941 g/mol = 4.93 Li and H2 ratio: 2:1 mol H2 = 4.93 / 2 = 2.47 V = nRT/P = 2.47 * 0.08206 * 299 K / 1.21 atm = 50.0 L

OpenStudy (anonymous):

Is my answer correct?

OpenStudy (anonymous):

@ganeshie8 @e.mccormick

OpenStudy (taramgrant0543664):

Yep that's right!

OpenStudy (anonymous):

Thanks! :)

OpenStudy (taramgrant0543664):

No problem!